Solve the following. Show your solution and box your final answer.
The OH- ion concentration of a blood sample is 2.5x10-7 M. What is the pH of the blood?
1
Expert's answer
2020-11-26T05:10:09-0500


You are given the OH concentration, so you can solve for the pOH.

"pOH + pH= 14\newline pOH= -log(OH)=-log(2.5×10^-7)=6.60\newline pH=14-pOH\newline =14-6.60\newline 7.40"